The length of a rectangle is 1 unit and the width is w.if the width is increased by 3 units,by how many units will the perimeter be increased?
Answer by fcabanski(1391)

P = 2L + 2W


For L = 1 that is:


P = 2+2W


If the width increases by 3 units it becomes w+3.
P=2(1) + 2(w+3) = 2 + 2W + 6 = 2W+8


It has increased by 6


2W + 8 - (2 + 2W) = 2W + 8 - 2 - 2W = 6

If you need help understanding math so you can solve these problems yourself, then one on one online tutoring is the answer ($30/hr). If you need faster solutions with guaranteed detailed answers, then go with personal problem solving ($3.50-$5.50 per problem). Contact me at fcabanski@hotmail.com